OpenStudy (anonymous):

Find \(x\). \[2^{2x} = 8^{x-1}\] \[8^{x-1}\Rightarrow (2^3)^{x-1}\] \[2^{2x} = (2^3)^{x-1}\] What do I do from here on?

terenzreignz (terenzreignz):

Apply the laws of exponents on the right-hand side \[\Large (a^n)^m = a^{mn}\]

OpenStudy (anonymous):

Ah, so I basically distribute? \[2^{2x} = (2^3)^{x-1}\] \[2^{2x} = 2^{3x-3}\] So from there: \[2x = 3x - 3\] \[-x = -3\] \[x = 3\] I have a question though, what if the bases were different?

terenzreignz (terenzreignz):

Like? (The bases WERE different, by the way)

terenzreignz (terenzreignz):

But strictly speaking, if you have something like this (I'm going to be very general here) \[\Large p^{f(x)}= q^{g(x)}\]

OpenStudy (anonymous):

Such as: \[2^{2x} = 4^{3x-3}\] Would that effect my answer in finding \(x\)?

terenzreignz (terenzreignz):

Well yes, but this case is actually rather mild, since you know fully well that \[\Large 2^2 = 4\implies=4^{3x-3}= 2^{2(3x-3)}\]

OpenStudy (anonymous):

What if it wasn't mild such as going to 5 or some other odd number?

terenzreignz (terenzreignz):

In general, when given \[\Large p^{f(x)}= q^{g(x)}\] The way to do it is to realise that \[\Huge q = p^{\log_p (q)}\] So that you get \[\Huge p^{f(x)}= p^{[g(x)]\log_p(q)}\] And finally \[\Large f(x) = g(x)\log_p(q) \]

terenzreignz (terenzreignz):

If we apply this to the first question, we, in fact, get \[\Large p = 2 \\\Large q=8\\\Large f(x) = 2x\\\Large g(x) = x-1\] Following that rule, we get \[\Large f(x) = g(x) \log_p(q)\]\[\Large 2x= (x-1)\log_2(8)\]\[\Large 2x = (x-1)\cdot 3 \] etc

terenzreignz (terenzreignz):

But this is 'mild' since 8 is a power of 2. The uglier solutions arise when q is not a power of p (or vice versa)

OpenStudy (anonymous):

I see, thank you.